# Towards High-Quality and Efficient Speech Bandwidth Extension with Parallel Amplitude and Phase Prediction

### v1

@@ V-A1 Baseline Methods
sinc 7.13 4.09 99.76 12.39 8.09 89.91 42.90 32.56 79.04
TFiLM [21] 7.22 4.20 99.24 15.16 10.51 91.27 55.97 42.59 80.23
AFiLM [22] 7.13 4.08 98.54 14.30 9.29 90.51 59.95 45.69 76.83
NVSR [31] 7.47 4.19 98.84 15.91 10.55 92.04 53.32 41.10 82.38
AERO [32] 7.17 4.08 99.38 11.75 7.39 93.74 - - -
AP-BWE 7.03 3.99 99.77 9.38 5.64 94.75 30.22 22.13 87.00
Ground Truth 6.13 3.32 100.00 6.13 3.32 100.00 6.13 3.32 100.00
V-A2 Evaluation on Speech Quality
As depicted in Table I, our proposed AP-BWE achieved the best performance in speech quality with all kinds of input sampling rates. Compared to sinc filter interpolation, our proposed AP-BWE exhibited significant improvements of 61.7%, 67.5%, and 68.6% in terms of LSD as well as 8.5%, 22.2%, and 37.7% in terms ViSQOL for input sampling rates of 8 kHz, 4 kHz, and 2 kHz, respectively. With the narrowing of input bandwidth, the performance advantage of our proposed AP-BWE became more pronounced, indicating the powerful BWE capability of our model. In general, waveform-based methods (TFiLM and AFiLM) performed less effectively than spectrum-based methods (NVSR, AERO, and our proposed AP-BWE), indicating the importance of capturing time-frequency domain characteristics for the BWE task. Within spectrum-based methods, NVSR, relying on high-frequency mel-spectrogram prediction and vocoder-based waveform reconstruction, demonstrated advantages in the LSD metric assessing the extended amplitude. However, the vocoder-based phase recovery was not as effective as the complex spectrum-based approach, so it lagged behind AERO in the ViSQOL metric assessing overall speech quality. Compared to AERO, our AP-BWE, benefiting from explicit amplitude and phase optimizations, successfully avoided the compensation effects between amplitude and phase and consequently achieved better performance in both spectral and waveform-based metrics. Worth noting is that the unified AP-BWE* model exhibited only a slight decrease in performance compared to AP-BWE, and even achieved the highest ViSQOL score at the input sampling rate of 2 kHz. This indicated that our model exhibited strong adaptability to the input sampling rate.
The key distinction between our approach and others lay in our implementation of explicit high-frequency phase extension. As illustrated in Table II, our proposed AP-BWE consistently outperformed other baselines across various input sampling rates, demonstrating superior performance in terms of instantaneous phase error and phase continuity along both time and frequency axes. For the AP-BWE*, only slight decreases in the AWPD metrics were observed when the input sampling rate was 8 kHz. Under other input sampling rate conditions, the metrics were the same as AP-BWE, indicating the robustness of our unified model on phase. Remarkably, for other baseline methods, some of their AWPD metrics exhibited degradation compared to those of the input sinc-interpolated waveforms. This suggests a limitation in the effective utilization of low-frequency phase information during the speech BWE process by these baseline methods. Moreover, all methods here directly generated waveforms without substituting the original low-frequency components, so their low-frequency phase might be partially compromised, leading to a significant impact on the quality of the extended speech. This observation underscored the critical importance of precise phase prediction and optimization in the context of BWE tasks, further emphasizing the advantage of our approach.
To more intuitively compare the BWE effects between our model and other baseline models, we visualized the spectrograms of the natural wideband 16 kHz speech waveform and speech waveforms extended by AP-BWE and other baseline methods at an input sampling rate of 2 kHz as illustrated in Fig. 4a. In general, TFiLM and AFiLM exhibited insufficient modeling capability for high-frequency components compared to NVSR and our proposed AP-BWE, particularly in the case of high-frequency unvoiced segments. Compared to NVSR, we can observe from the spectrogram that our result had higher energy in both high-frequency unvoiced segments (e.g., 0.2 $\sim$ 0.3s) and low-frequency harmonics (e.g., 1.1 $\sim$ 1.5s). As a result, the speech signals extended by our proposed AP-BWE would sound brighter, contributing to a better BWE effect. Compared to the natural wideband speech, even in cases where the input bandwidth was as narrow as 1 kHz, our model was still capable of extending a relatively complete harmonic structure. This highlighted the robust modeling capability of our proposed AP-BWE and the effectiveness of explicit predictions of amplitude and phase spectra.
V-A3 Evaluation on Generation Efficiency
We respectively evaluated the generation efficiency of our proposed AP-BWE as well as other baseline methods on both CPU and GPU as outlined in Table I. Since NVSR divided the BWE process into the mel-spectrogram extension stage and vocoder synthesis stage, it lagged far behind other end-to-end methods in terms of inference speed. For TFiLM and AFiLM, since they both operated on the waveform level and utilized RNNs or self-attention to capture long-term dependencies, their generation efficiencies were consequently constrained. For AERO, although it and our proposed AP-BWE both operated on the spectral level, the utilization of transformer [43] blocks in multiple layers severely slowed down its inference speed. Nevertheless, our AP-BWE model, based on fully convolutional networks and all-frame-level operations, has achieved an astonishingly high-speed waveform generation (29.61 times real-time speed on CPU and 382.56 times on GPU), far surpassing other baseline methods.
@@ V-A4 Evaluation on Speech Intelligibility
As shown in Table III, it is obvious that our proposed AP-BWE exhibited a remarkable improvement in terms of intelligibility metrics compared to baseline models. Although all baseline models showed some improvement in the STOI metric under every input sampling rate, when it comes to the WER and CER metrics, except for AERO and our proposed AP-BWE, other baseline methods exhibited an increase in both WER and CER compared to sinc filter interpolation, indicating that frequency bands they extend, on the contrary, reduced the intelligibility of the waveforms. It is noteworthy that, under the condition of extending from 8 kHz to 16 kHz, the intelligibility metrics of sinc filter interpolation were already very close to the Ground Truth. However, when the input sampling rate turned to 2 kHz, with a bandwidth of only 1 kHz, the intelligibility metrics sharply deteriorated. Overall, these baseline models demonstrated limited extension capabilities in scenarios with extremely constrained input frequency bands. In contrast, at the input sampling rate of 2 kHz, our proposed AP-BWE has still shown improvements of 29.6%, 32.0%, and 10.1% in WER, CER, and STOI on the basis of sinc filter interpolation. This indicated that benefiting from our precise phase prediction, our model possessed strong harmonic restoration capabilities, reconstructing the key information of vowels and consonants as well as significantly enhancing the intelligibility of the extended speech.
TABLE IV: Experimental Results in Speech Quality (LSD and ViSQOL) and Generation Efficiency (RTF) for BWE Methods Evaluated on the VCTK Dataset with Target Sampling Rate of 48 kHz, Where in RTF ($a\times$) Representing $a$ Times Real-Time
Method $\mathrm{24kHz\rightarrow 48kHz}$ $\mathrm{16kHz\rightarrow 48kHz}$ $\mathrm{12kHz\rightarrow 48kHz}$ $\mathrm{8kHz\rightarrow 48kHz}$ $\mathrm{RTF(CPU)}$ $\mathrm{RTF(GPU)}$
@@
sinc 2.17 2.99 2.57 2.26 2.75 2.09 2.94 2.07 - -
NU-Wave [46] 0.85 3.18 0.99 2.36 - - - - 95.57 (0.01$\times$) 0.5018 (1.99$\times$)
NU-Wave2 [47] 0.72 3.74 0.86 3.00 0.94 2.75 1.09 2.48 92.58 (0.01$\times$) 0.5195 (1.92$\times$)
UDM+ [48] 0.64 4.02 0.79 3.55 0.88 3.08 1.03 2.81 74.03 (0.01$\times$) 0.8335 (1.20$\times$)
mdctGAN [33] 0.72 3.58 0.82 3.15 0.90 2.96 0.93 2.95 0.2461 (4.06$\times$) 0.0129 (77.80$\times$)
AP-BWE* 0.62 4.17 0.72 3.63 0.79 3.46 0.85 3.32 0.0551 (18.14$\times$) 0.0034 (292.28$\times$)
@@
TABLE V: Experimental Results for the Band-wise Analysis with Source Sampling Rate of 8 kHz and Target Sampling Rate of 48 kHz
@@ V-B1 Baseline Methods
For BWE targeting a 48 kHz sampling rate, the sinc filter interpolation was still used as the low-bound method. We subsequently compared our proposed AP-BWE with three diffusion-based methods (NU-Wave [46], NU-Wave 2 [47], and UDM+ [48]) and an MDCT-spectrum-based method (mdctGAN [33]). For NU-Wave, we used the community-contributed checkpoints from their official implementation  [^8] https://github.com/maum-ai/nuwave.. Notably, NU-Wave did not conduct the experiment at input sampling rates of 8 kHz and 12 kHz, so we excluded these results from our analysis. For NU-Wave 2 and UDM+, we used the reproduced NU-Wave 2 checkpoint and official UDM+ checkpoint  [^9] https://github.com/yoyololicon/diffwave-sr.. For mdctGAN, It is worth noting that in the original paper of mdctGAN, the mdctGAN model was trained on the combination of the VCTK training set and the HiFi-TTS dataset, and tested on the VCTK test set. Here, we employed the checkpoints from its official implementation  [^10] https://github.com/neoncloud/mdctGAN.. In addition, the AP-BWE* was trained using randomly selected sampling rates from 8 kHz, 12 kHz, 16 kHz, and 24 kHz to handle inputs of various resolutions.
(a) 2 kHz $\rightarrow$ 16 kHz
(b) 8 kHz $\rightarrow$ 48 kHz
Figure 4: Spectrogram visualization of the original wideband speech waveform and speech waveforms extended by baseline methods and our proposed AP-BWE.
@@ V-B2 Evaluation on Speech Quality
As depicted in Table IV, for the high-sampling-rate waveform generation at 48 kHz, our proposed AP-BWE still achieved the SOTA performance, irrespective of the input sampling rates. In general, compared to the baseline models, our approach exhibited a notably significant improvement in ViSQOL, particularly at lower input sampling rates, which underscored the substantial impact of precise phase prediction on speech quality. For diffusion-based methods, since both NU-Wave2 and UDM+ implemented a single model for extension across different input sampling rates, we compared our unified AP-BWE* model with them. Compared to NU-Wave 2 and UDM+, our AP-BWE* model exhibited growing superiority in LSD as the input sampling rate decreased. This suggested that diffusion-based methods, operating at the waveform level, struggled to effectively recover spectral information in scenarios with restricted bandwidth. Although both mdctGAN and our proposed AP-BWE were spectrum-based methods, AP-BWE significantly outperformed it across all input sampling rates. Especially in terms of the overall speech quality, our proposed AP-BWE surpassed mdctGAN by 18.7%, 17.5%, 16.9%, and 13.6% in ViSQOL at input sampling rates of 24 kHz, 16 kHz, 12 kHz, and 8 kHz, respectively. This suggested that STFT spectra were more suitable for waveform generation tasks compared to MDCT spectra. Additionally, similar to the results at 16 kHz, our unified AP-BWE* model demonstrated competence across different sampling rate inputs, with only a slight decline in quality compared to AP-BWE, reaffirming the adaptability of our approach to input sampling rates.
@@
Unlike the strong harmonic structure observed in the high-frequency components of 16 kHz waveforms, the high-frequency portion of 48 kHz waveforms exhibited more randomness. In our preliminary experiments, we observed the phase metrics of extended 48 kHz waveforms show minimal variation between systems, especially in scenarios with relatively higher input sampling rates. Therefore, with the input sampling rate of 8 kHz, we calculated LSD and AWPD separately for different frequency bands of the extended 48 kHz waveform to assess the performance of these models within different frequency ranges, and the evaluation results are depicted in Table V. For the LSD metric, our AP-BWE outperformed other baseline models within each frequency band. Regarding the AWPD metrics, we only exhibited an advantage in the 4 kHz $\sim$ 8 kHz frequency band, while the differences between systems were minimal in the 8 kHz $\sim$ 12 kHz and 12 kHz $\sim$ 24 kHz frequency bands. This indicated that our proposed AP-BWE, benefiting from explicit phase prediction, was capable of effectively recovering the harmonic structure in the waveform, thereby significantly improving the speech quality in the mid-to-low-frequency range. For high-frequency phases, due to their strong randomness, the current methods exhibited comparable predictive capabilities.
@@
We also visualized the spectrograms of the wideband 48 kHz speech waveform and speech waveforms extended by AP-BWE and other baseline methods at an input sampling rate of 8 kHz as illustrated in Fig. 4b. The spectrograms clearly revealed that NU-Wave2 and UDM+ displayed relatively weaker recovery capabilities for high-frequency components, particularly with low energy in the unvoiced segments, resulting in extended speech that sounded less bright. This finding aligned with the results obtained at the target sampling rate of 16 kHz, suggesting a potential limitation in modeling high-frequency unvoiced segments for waveform-based methods. Comparing our proposed AP-BWE with mdctGAN, it can be observed that the spectrogram generated by mdctGAN exhibited higher randomness and poorer continuity, resulting in a less stable auditory perception. In contrast, our proposed AP-BWE demonstrated robust restoration capabilities for the high-frequency components, especially in the unvoiced segments. While there were still differences in the high-frequency components of the voiced segments compared to natural wideband speech, these distinctions had minimal impact on the perceptual quality of the speech.
@@ V-B3 Evaluation on Generation Efficiency
As depicted in Table IV, our proposed AP-BWE remained capable of efficient 48 kHz speech waveform production at a speed of 18.14 times real-time on CPU and 292.28 times real-time on GPU. For diffusion-based methods (i.e., NU-Wave, NU-Wave2, and UDM+), their generation efficiency took a significant hit as they require multiple time steps in the reverse process to continuously denoise and recover the extended waveform from latent variables. Remarkably, our AP-BWE achieved a speedup over them by approximately 1000 times on CPU and 100 times on GPU, respectively. Despite both mdctGAN and our proposed AP-BWE operating at the spectral level, the generation speed of mdctGAN was still constrained by its two-dimensional convolution and transformer-based structure. Consequently, our AP-BWE which is fully based on one-dimensional convolutions enabled an approximately fourfold acceleration compared to it. Compared to running on a GPU, our model exhibited a more significant efficiency improvement on the CPU. This indicated that our model can efficiently generate high-sampling-rate samples even without the parallel acceleration support of GPUs, making it more suitable for applications in scenarios with limited computational resources. Comparing Table I and Table IV, it can be observed that, for generating speech waveforms of the same duration, the inference efficiency at a 48 kHz sampling rate was relatively lower than that at 16 kHz. This was attributed to the fact that our model under both sampling rate configurations utilized the same STFT settings, resulting in different frame numbers processed by the model.
Figure 5: Spectrogram visualization of the original wideband speech waveform and speech waveforms extended by the ablation models of our proposed AP-BWE with an input sampling rate of 8 kHz and target sampling rate of 48 kHz. “AP-BWE w/o MRDs” represents the ablation of both MRAD and MRPD, while “AP-BWE w/o Disc.” denotes the ablation of all discriminators.

### v2

@@ V-A3 Evaluation on Generation Efficiency
sinc 3.67 1.67 99.76 11.45 7.08 89.91 47.43 33.56 79.04
TFiLM [21] 3.69 1.69 99.24 11.32 7.24 91.27 45.95 33.58 80.23
AFiLM [22] 3.67 1.67 98.54 9.28 5.53 90.51 45.16 33.01 76.83
NVSR [31] 4.38 2.02 98.84 13.56 8.51 92.04 59.53 44.43 82.38
AERO [32] 3.97 1.84 99.38 9.78 5.51 93.74 - - -
AP-BWE 3.72 1.67 99.77 6.69 3.54 94.75 36.69 25.61 87.00
Ground Truth 3.07 1.26 100.00 3.07 1.26 100.00 3.07 1.26 100.00
@@ V-A4 Evaluation on Speech Intelligibility
As shown in Table IV, it is obvious that our proposed AP-BWE exhibited a remarkable improvement in terms of intelligibility metrics compared to baseline models. Under the condition of extending from 8 kHz to 16 kHz, the performance of the sinc filter interpolation was already very close to the Ground Truth. Our proposed AP-BWE and other baseline models struggled to further improve WER and CER on top of the waveform interpolated by the sinc filter, suggesting that the ASR model focused on information from frequencies below 4 kHz for transcription. When the source sampling rate was further reduced to 4 kHz and 2 kHz, all the baseline models showed slight improvements in WER and CER compared to sinc filter interpolation, except for NVSR. The decline in NVSR’s performance in WER and CER was due to its use of a vocoder to restore the waveform, which made the low-frequency components unnatural, but its STOI metric was still improved. Overall, these baseline models demonstrated limited extension capabilities under the extremely high extension ratio. However, our proposed AP-BWE significantly improved WER, CER, and STOI by 41.57%, 50.00%, 5.38% at the 4 kHz source sampling rate, and by 22.64%, 23.69%, and 10.07% at the 2 kHz source sampling rate, compared to sinc filter interpolation. This indicated that benefiting from our precise phase prediction, our model possessed strong harmonic restoration capabilities, reconstructing the key information of vowels and consonants as well as significantly enhancing the intelligibility of the extended speech.
TABLE V: Experimental Results in Speech Quality (LSD and ViSQOL) and Generation Efficiency (RTF and FLOPs) for BWE Methods Evaluated on the VCTK Dataset with Target Sampling Rate of 48 kHz, Where in RTF ($a\times$) Representing $a$ Times Real-Time
Method $\mathrm{24kHz\rightarrow 48kHz}$ $\mathrm{16kHz\rightarrow 48kHz}$ $\mathrm{12kHz\rightarrow 48kHz}$ $\mathrm{8kHz\rightarrow 48kHz}$ $\mathrm{RTF(CPU)}$ $\mathrm{RTF(GPU)}$ $\mathrm{FLOPs}$
@@
sinc 2.17 2.99 2.57 2.26 2.75 2.09 2.94 2.07 - - -
NU-Wave [46] 0.85 3.18 0.99 2.36 - - - - 95.57 (0.01$\times$) 0.5018 (1.99$\times$) 4039.13G
NU-Wave2 [47] 0.72 3.74 0.86 3.00 0.94 2.75 1.09 2.48 92.58 (0.01$\times$) 0.5195 (1.92$\times$) 1385.27G
UDM+ [48] 0.64 4.02 0.79 3.35 0.88 3.08 1.03 2.81 74.03 (0.01$\times$) 0.8335 (1.20$\times$) 2369.50G
mdctGAN [33] 0.71 3.69 0.83 3.27 0.85 3.12 0.93 3.03 0.2461 (4.06$\times$) 0.0129 (77.80$\times$) 103.38G
AP-BWE* 0.62 4.17 0.72 3.63 0.79 3.46 0.85 3.32 0.0551 (18.14$\times$) 0.0034 (292.28$\times$) 17.87G
@@
TABLE VI: Experimental Results for the Band-wise Analysis with Source Sampling Rate of 8 kHz and Target Sampling Rate of 48 kHz
@@ V-B1 Baseline Methods
For BWE targeting a 48 kHz sampling rate, the sinc filter interpolation was still used as the low-bound method. We subsequently compared our proposed AP-BWE with three diffusion-based methods (NU-Wave [46], NU-Wave 2 [47], and UDM+ [48]) and an MDCT-spectrum-based method (mdctGAN [33]). For NU-Wave, we used the community-contributed checkpoints from their official implementation  [^7] https://github.com/maum-ai/nuwave.. Notably, NU-Wave did not conduct the experiment at source sampling rates of 8 kHz and 12 kHz, so we excluded these results from our analysis. For NU-Wave 2 and UDM+, we used the reproduced NU-Wave 2 checkpoint and official UDM+ checkpoint  [^8] https://github.com/yoyololicon/diffwave-sr.. It is worth noting that in the original paper of mdctGAN [33], the mdctGAN model was trained on the combination of the VCTK training set and the HiFi-TTS dataset, and tested on the VCTK test set. Here, for a fair comparison, we re-trained all the mdctGAN models solely on the VCTK training set following its official implementation  [^9] https://github.com/neoncloud/mdctGAN.. In addition, the AP-BWE* was trained using randomly selected sampling rates from 8 kHz, 12 kHz, 16 kHz, and 24 kHz to handle inputs of various resolutions.
@@ V-B2 Evaluation on Speech Quality
    As depicted in Table V, for the high-sampling-rate waveform generation at 48 kHz, our proposed AP-BWE still achieved the SOTA performance in objective metrics, irrespective of the source sampling rates. In general, compared to the baseline models, our approach exhibited a notably significant improvement in ViSQOL, particularly under lower extension ratios, which underscored the substantial impact of precise phase prediction on speech quality. For diffusion-based methods, since both NU-Wave2 and UDM+ implemented a single model for extension across different source sampling rates, we compared our unified AP-BWE* model with them. Compared to NU-Wave 2 and UDM+, our AP-BWE* model exhibited growing superiority in LSD as the source sampling rate decreased. This suggested that diffusion-based methods, operating at the waveform level, struggled to effectively recover spectral information in scenarios with restricted bandwidth. Although both mdctGAN and our proposed AP-BWE were spectrum-based methods, AP-BWE significantly outperformed it across all source sampling rates. Especially in terms of the overall speech quality, our proposed AP-BWE surpassed mdctGAN by 15.2%, 13.1%, 10.9%, and 10.6% in ViSQOL at source sampling rates of 24 kHz, 16 kHz, 12 kHz, and 8 kHz, respectively. This suggested that STFT spectra were more suitable for waveform generation tasks compared to MDCT spectra. Additionally, similar to the results at 16 kHz, our unified AP-BWE* model demonstrated competence across different sampling rate inputs, with only a slight decline in quality compared to AP-BWE, reaffirming the adaptability of our approach to source sampling rates.
@@
    Unlike the strong harmonic structure observed in the high-frequency components of 16 kHz waveforms, the high-frequency portion of 48 kHz waveforms exhibited more randomness. In our preliminary experiments, we observed the phase metrics of extended 48 kHz waveforms show minimal variation between systems, especially in scenarios with relatively higher source sampling rates. Therefore, with the source sampling rate of 8 kHz, we calculated LSD and AWPD separately for different frequency bands of the extended 48 kHz waveform to assess the performance of these models within different frequency ranges, and the evaluation results are depicted in Table VI. For the LSD metric, our AP-BWE outperformed other baseline models within each frequency band. Regarding the AWPD metrics, we only exhibited an advantage in the 4 kHz $\sim$ 8 kHz frequency band, while the differences between systems were minimal in the 8 kHz $\sim$ 12 kHz and 12 kHz $\sim$ 24 kHz frequency bands. This indicated that our proposed AP-BWE, benefiting from explicit phase prediction, was capable of effectively recovering the harmonic structure in the waveform, thereby significantly improving the speech quality in the mid-to-low-frequency range. For high-frequency phases, due to their strong randomness, the current methods exhibited comparable predictive capabilities.
    Figure 5: Spectrogram visualization of the original wideband 48 kHz speech waveform and speech waveforms extended by baseline methods and our proposed AP-BWE from the source sampling rate of 8 kHz.
    TABLE VII: MOS Tests Results for BWE Methods with Source Sampling Rate of 8 kHz and Target Sampling Rate of 48 kHz
    Methods MOS (CI)
    sinc 3.69 ($\pm$ 0.07)
    NU-Wave2 [47] 3.75 ($\pm$ 0.08)
    UDM+ [48] 3.98 ($\pm$ 0.06)
    mdctGAN [33] 4.00 ($\pm$ 0.07)
    AP-BWE 4.11 ($\pm$ 0.06)
    Ground Truth 4.17 ($\pm$ 0.06)
@@
  • Subjective Evaluation: As shown in the right half of Table VII, we conducted MOS tests on the wideband 48 kHz speech waveform, along with speech waveforms extended by AP-BWE and other baseline methods at a source sampling rate of 8 kHz. We also visualized the corresponding spectrograms, as illustrated in Fig. 5. In this configuration, the initial 4 kHz bandwidth already contained the full fundamental frequency and most harmonic structures, resulting in less pronounced subjective listening differences between the speech extended by different models. Nevertheless, the MOS results demonstrated that our proposed AP-BWE still showed substantial advantages in subjective quality over other baseline models ($p<0.05$). Firstly, NU-Wave2 scored very significantly lower in MOS compared to our proposed AP-BWE ($p<0.01$), showing only a slight improvement over sinc filter interpolation, with spectrogram analysis revealing poor recovery of mid-to-high frequency components. UDM+ performed well in recovering mid-frequency components of speech, but it seemed to struggle with restoring higher-frequency components, particularly with low energy in the unvoiced segments, resulting in extended speech that sounded less bright. Consequently, the subjective quality of UDM+ remained significantly lower than that of our proposed AP-BWE ($p=0.020$). This finding aligned with the results obtained at the target sampling rate of 16 kHz, suggesting a potential limitation in modeling high-frequency unvoiced segments for waveform-based methods. The mdctGAN achieved the optimal MOS among the baseline methods, with its corresponding spectrogram displaying brighter and more complete structures. However, the high-frequency components of the spectrogram exhibited higher randomness and poorer continuity, resulting in a less stable auditory perception. In contrast, our proposed AP-BWE demonstrated a more robust restoration capability for the high-frequency components, especially in the unvoiced segments, giving it a substantial advantage in subjective speech quality over mdctGAN ($p=0.026$). While there were still differences in the high-frequency components of the voiced segments compared to natural wideband speech, these distinctions had minimal impact on the perceptual quality of the speech. Therefore, AP-BWE achieved a MOS close to that of natural wideband speech.
@@ V-B3 Evaluation on Generation Efficiency
Considering the inference speed, as depicted in Table V, our proposed AP-BWE remained capable of efficient 48 kHz speech waveform production at a speed of 18.14 times real-time on CPU and 292.28 times real-time on GPU. For diffusion-based methods (i.e., NU-Wave, NU-Wave2, and UDM+), their generation efficiency took a significant hit as they require multiple time steps in the reverse process to continuously denoise and recover the extended waveform from latent variables. Remarkably, our AP-BWE achieved a speedup over them by approximately 1000 times on CPU and 100 times on GPU, respectively. Despite both mdctGAN and our proposed AP-BWE operating at the spectral level, the generation speed of mdctGAN was still constrained by its two-dimensional convolution and transformer-based structure. Consequently, our AP-BWE which is fully based on one-dimensional convolutions enabled an approximately fourfold acceleration compared to it. Compared to running on a GPU, our model exhibited a more significant efficiency improvement on the CPU. This indicated that our model could efficiently generate high-sampling-rate samples even without the parallel acceleration support of GPUs, making it more suitable for applications in scenarios with limited computational resources. Considering the models’ computational complexity, the FLOPs of diffusion models are heavily constrained by their reverse steps (82.78G $\times$ 50 steps for NU-Wave, 27.71G $\times$ 50 steps for NU-Wave2, and 47.39G $\times$ 50 steps for UDM+). However, even with just one step generation, the FLOPs of our proposed AP-BWE were still smaller than theirs, further demonstrating the superiority of our proposed AP-BWE in terms of generation efficiency. Comparing Table I and Table V, it can be observed that for generating speech waveforms of the same duration, the inference speed of our model at 48 kHz sampling rate was relatively lower, while the computational complexity was higher compared to the 16 kHz sampling rate. This was attributed to the fact that our model under both sampling rate configurations utilized the same STFT settings, resulting in different frame numbers processed by the model.
Figure 6: Spectrogram visualization of the original wideband speech waveform and speech waveforms extended by the ablation models of our proposed AP-BWE with a source sampling rate of 8 kHz and target sampling rate of 48 kHz. “AP-BWE w/o MRDs” represents the ablation of both MRAD and MRPD, while “AP-BWE w/o Disc.” denotes the ablation of all discriminators.